Based on your initial research, you found out that the human body can survive an acceleration trauma incident (sudden stop) if the magnitude of the acceleration is less than 250m/s2.what must be the inflation distance of airbag from car dashboard to avoid head injuries?
1
Expert's answer
2020-09-28T08:05:32-0400

Assume that we have initial velocity "v" and acceleration required to save human's live "a=250\text{ m/s}^2." The distance that will provide such acceleration is


"d=\frac{v^2}{2a}=\frac{v^2}{500}."

For example, for 60 km/h this will be 55 cm.